[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Sep 15, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Sep 16 03:31:54 UTC 2016


HSU Shift Log for Thu Sep 15 2016 23:31:48 GMT-0400 (EDT)

Shift: R

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: ATCF did not correctly repopulate the ATCF forecast track  
spread with for TS Ian with new forecast positions after late-arriving  
ASCAT-B pass showed center was at least 60 nmi nw of 00Z satellite  
estimates; tried Adv. Comp. twice, but had to finally close out and re-open  
ATCF window, which cleared up problem. TCD on TS Ian was late as a result.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160916/45fb2cfc/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list